13)Show that the function f (z)=(z̅ )²/z,z≠0 ;0, z=0 satiesfies Cauchy-Rieman equations at z=0.Does f'(0) exist? Explain the problem with step by step process?


1
Expert's answer
2021-09-24T02:14:28-0400

Consider

"f(z)= |z|^2\\
det \space \Delta w = f(z + \Delta z)- f(z)\\
\frac{ \Delta w}{ \Delta z}=\frac{ |z+\Delta z|^2-|z|^2}{ \Delta z}\\
\frac{ \Delta w}{ \Delta z}=\frac{ (z+\Delta z)(\bar{z}+\bar{\Delta z})-z \bar{z}}{ \Delta z}\\
\frac{ \Delta w}{ \Delta z}=\bar{z}+\bar{\Delta z}+\frac{ z\bar{\Delta z}}{ \Delta z}\\"

When the horizontal line and vertical line approach "\bar{\Delta z}" towards the origin

"\bar{\Delta z}={\Delta z}, \bar{\Delta z}=-{\Delta z}"

Then

"\frac{\Delta w}{\Delta z}=\bar z+\Delta z+z" for horizontal line


"\frac{\Delta w}{\Delta z}=\bar z-\Delta z-z" for vertical line

as "\Delta z \to 0"

"\bar z+\Delta z=\bar z-\bar z" or z= 0

So f'(0) does not exist when "z \not=0"